# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

### v1

@@ 4.3. Convergence rates using a-posteriori stopping criteria
Proof.
We apply (4.30) along with (4.33) with $k=k_{\delta}$ to obtain
$k_{\delta}^{\gamma+1/2}\leq c_{\tau\eta}C_{\gamma}\|z\|/\delta$
with $C_{\gamma}$ as in Lemma 4.4. Hence, we obtain
$k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
where $C_{\gamma,\eta,n,\tau,\theta}=(c_{\tau\eta}C\gamma)^{\frac{2}{2\gamma+1}}.$ The error estimate (4.40) originates from (4.25). ∎
4.4. Convergence rates using heuristic type stopping criteria
In this section, we will derive an error estimate for the expression $u^{\delta}_{k_{*}(y^{\delta})}-u^{\dagger}$ in mean squared norm.
Lemma 4.5.
Let Assumption 3.1 hold and $y^{\delta}$ be a noisy data satisfying Assumption 3.2, also let $k_{*}:=k_{*}(y^{\delta})$ be the integer defined by Rule 1. Additionally, assume that
((4.41)) $\|F(u_{0})-y^{\delta}\|>\tau\|y^{\dagger}-y^{\delta}\|.$
If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
((4.42)) $(k_{*}+a)\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
where $C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}=\frac{\tau^{2}(1+a)C_{\gamma,\eta,n,\tau,\theta}}{\varkappa^{2}}$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
Proof.
By Assumption 3.2 and the definition of $k_{*}$ and $k_{\delta}$, we have
$\varkappa^{2}(k_{*}+a)\delta^{2}\leq\Psi(k_{*},y^{\delta})\leq\Psi(k_{\delta},y^{\delta})=(k_{\delta}+a)\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|^{2}\leq\tau^{2}(k_{\delta}+a)\delta^{2}.$
This implies that
$\varkappa^{2}(k_{*}+a)\leq\tau^{2}(k_{\delta}+a)\leq\tau^{2}(1+a)k_{\delta},\ \text{as}\ k_{\delta}\geq 1$
since $\|F(u_{0})-y^{\delta}\|>\tau\delta$. Therefore, by virtue of Theorem 4.3, the proof is complete. ∎
Lemma 4.6.
Let Assumption 3.1 hold and let $k_{*}$ be the integer defined by Rule 1 and let $y^{\delta}$ satisfy (1.3). If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
((4.43)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma},\ \ \ \text{and}\ \ \ \mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma-1/2}$
for $0\leq k\leq k_{*},$ $\tilde{C}_{\gamma}>0$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
Proof.
The proof follows on the similar lines as given in Lemma 4.4 with a small modification in (4.30) due to heuristic type stopping rule $k_{*}$. By virtue of Lemma 4.5, we have
((4.44)) $\delta\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|(k_{*}+a)^{-\gamma-1/2}\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|j^{-\gamma-1/2}.$
Using (4.44) in (4.28) and (4.29) and following the same procedure as in Lemma 4.4 we get the desired result. ∎
Remark 4.3.
In comparison to Lemma 4.4, the only difference in Lemma 4.6 is a constant factor. This variation arises due to (4.44), which is applied in the proof in place of (4.30).
Theorem 4.4.
Under the assumptions of Lemma 4.6 we have
((4.45)) $\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}},$
with some constants $\tilde{C}_{\gamma},C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}>0,$ depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
Proof.
We apply (4.42) along with (4.43) with $k=k_{*}$ as
$\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}$ $\leq\tilde{C}_{\gamma}\|z\|C^{-\gamma}_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\left(\frac{\|z\|}{\delta}\right)^{\frac{-2\gamma}{2\gamma+1}}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}}.$
Hence we get the desired result.
4.5. Comparison of convergence rates
In this subsection, we present a detailed comparison of various iterative methods. As shown in Table 1, the convergence rates achieved by our approach are comparable to those reported in the existing literature. Furthermore, we selected the step size $\omega_{k}\in[\omega,\Omega]$ in such a way that it can either remain constant or decay monotonically, depending on the specific method used.
Method Step-size Stopping criteria Convergence Rate
LM [8] Constant a posteriori (1.7) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma>0$
IRLM [16] Constant a-priori $O(\delta^{1/2})$
SGD [14] Decaying $k^{*}=\left[(\frac{\delta}{\|w\|})^{-\frac{2}{(2\gamma+1)(1-\alpha)}}\right]$ $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$ and $\alpha$ close to $1$
SGD [12] Decaying a posteriori (1.7) Not available
IRSGD (This paper) Any a posteriori (1.7) $O(\delta^{\frac{2\gamma}{(2\gamma+1)}}),\gamma\in(0,\frac{1}{2})$
IRSGD (This paper) Any heuristic type (Rule 1) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$
Table 1. Convergence rates of various iterative methods.
5. Numerical Experiments
The objective of this section is to discuss the applicability of proposed scheme using Rule 1 on linear as well as nonlinear ill-posed inverse problems.
@@ 5.1. Linear Case
Consider the following problem of linear ill-posed system (see [15] for details on this problem)
[…]
derived from first-kind linear integral equations on $[a,b]$ by sampling at $t_{i}\in[a,b]$ for $i=0,\ldots,P-1,$ where $t_{i}=a+\frac{i(b-a)}{P-1}$ for $i=0,\ldots,P-1,$ and the kernel $k(\cdot,\cdot)$ is continuous on $[a,b]\times[a,b]$. In this case, the iterative scheme (3.1) becomes
[…]
where $i_{k}$ is the randomly chosen index at the $k$th iteration step from the index set $\{0,\ldots,P-1\}$. We consider the system (5.1) with $[a,b]=[-6,6],P=1000,$ and $k(t,s)=\varphi(t-s),$ where $\varphi(t)=(1+\cos(\frac{\pi t}{3}))_{\chi_{\{|t|<3\}}}$ to test the effectiveness of technique (5.2). We consider the following solution to be reconstructed
@@
Instead of the exact data $y_{i}=F_{i}(u^{\dagger})$, we use the noisy data
((5.3)) $y^{\delta}_{i}=y_{i}+\delta_{\text{rel}}\times|y_{i}|\times{\epsilon_{i}},\quad i=0,\ldots,P-1,$
where $\epsilon_{i}$ is a random noise that follows a standard Gaussian distribution $N(0,1)$, and $\delta_{\text{rel}}$ is the relative noise level. Based on the division of $[-6,6]$ into $P-1$ subintervals of similar length, the trapezoidal rule approximates the integrals involved in the procedure and discretized into a linear system of size $n=1000$ with the forward operator $F_{n\times n}$. We execute the method (5.2) with the numerical settings as follows:
@@
    The maximum number of iterations is set to $k_{\text{max}}=1000$.
@@
    We employ noisy data, where $\delta_{\text{rel}}=10^{-1},10^{-2},10^{-3}$ and $10^{-4}$ represent four different relative noise levels.
@@ 5.1.1. Convergence behavior of $\Psi(k,y^{\delta})$
Figure 1. Convergence behavior of $\Psi(k,y^{\delta})$ as defined in Rule 1, alongside the increase in the stopping index $k_{*}$ as the noise level decreases.
Figure 1 clearly shows that the noise level in the input data has a significant effect on the convergence behavior of $\Psi(k,y^{\delta})$. As $\delta_{\text{rel}}$ decreases from $10^{-1}$ to $10^{-4}$, convergence becomes both faster and more stable. Additionally, it demonstrates that $\Psi(k_{*},y^{\delta})\to 0$ as $\delta\to 0$, which supports the assertion in Lemma 3.2. Furthermore, it is also evident that as the noise level decreases, the stopping index $k_{*}$ increases.
@@ 5.1.2. Reconstructed solutions
Table 2. Results of IRSGD with different values of $a$ in Rule 1.
@@ 5.1.3. Comparison between SGD and IRSGD
We compare the results of IRSGD with SGD. For this comparison, we set the constant $a$ in the rule (referred to as Rule 1) to $a=100$. The numerical results for the comparison are presented in Table 3.
@@
Table 3. Comparison of IRSGD and SGD with Rule 1.
@@ 5.2. Nonlinear case
For each $i$ between 00 and $P-1$, let $\sigma_{i}\in S^{1}$ be a set of recording directions such that $\zeta_{i}=\zeta(\theta_{i})=(\cos\theta_{i},\sin\theta_{i})$. The problem is to reconstruct a function $u$ supported on a bounded domain $D:=[-1,1]\times[-1,1]\subset\mathbb{R}^{2}$ from the system of operator equations
$F_{i}(u)=y_{i}^{\delta},\quad\text{for }i=0,\ldots,P-1,$
@@
[…]
From [10], we know that each operator $F_{i}:H_{0}^{1}(D)\to L^{2}(I)$ with $I=[-1,1]$ is continuous and Fréchet differentiable with
$F_{i}^{\prime}(u)h=2R_{i}(u)R_{i}(h),\quad\forall h\in H_{0}^{1}(D).$
In addition, the adjoint $F_{i}^{\prime}(u)^{*}:L^{2}(I)\to H_{1}^{0}(D)$ is given by
@@
where $R_{i}^{*}:L^{2}(I)\to L^{2}(D)$ is the adjoint of $R_{i}$ given by $(R_{i}^{*}g)(u)=g(u\cdot\sigma_{i}),$ $I$ refers to the identity operator and $\Delta$ represents the Laplace operator acting on $H_{0}^{1}(D)$. The computation of the adjoint operator requires solving an elliptic boundary value problem, which is approximated using a multigrid method implemented through finite difference discretization.
In this numerical experiment, we set the parameters as follows: $\delta_{\text{rel}}=10^{-2}$, $k_{\text{max}}=750$, $\omega_{k}=\omega=5\times 10^{-3}$, and $\lambda_{k}=\frac{1}{k^{3}}$. The true image and initial guess are shown in Figure LABEL:initial_and_true_figures. Using the proposed method, we reconstructed the image with a stopping criterion based on Rule 1. We evaluated the reconstruction quality by tracking both the relative errors and the progression of $\Psi(k,y^{\delta})$ for different values of $a$: specifically, $a=10^{3}$ and $10^{4}$. The results for each case are displayed in Figures LABEL:RE=1000-LABEL:RE=_10000, respectively.
@@ 6. Conclusion
In this paper, we present an Iteratively Regularized Stochastic Gradient Descent (IRSGD) method specifically developed to address systems of ill-posed inverse problems, utilizing the Hanke-Raus heuristic rule. Notably, the proposed IRSGD method is fully data-driven, as the Hanke-Raus rule operates independently of any prior knowledge of the noise level. Our convergence analysis is based on the tangential cone condition and assumes Lipschitz continuity of the operators involved.
Furthermore, we have conducted a thorough analysis of the convergence of our method in the mean squared norm. Additionally, we have established convergence rates under both a-posteriori and heuristic rules by leveraging the source condition, range invariance condition, and stochastic range invariance condition for arbitrary step sizes. To demonstrate the practical effectiveness of our approach, we have provided extensive numerical experiments on both linear and nonlinear inverse problems.
Lastly, we outline several directions for future research. First, it is essential to validate our assumptions and theoretical analysis in the context of specific nonlinear inverse problems to further demonstrate the practical applicability of the IRSGD method. Second, following the work in [2, 4, 5, 19], the IRSGD approach can be extended to handle nonsmooth nonlinear problems, thereby broadening its range of applications. Finally, it is important to investigate adaptive step size rules to provide practical insights and guidelines for real-world implementation. The heuristic rule (based on the modified discrepancy principle), which is used in this analysis can be further modified to study the stochastic analysis of regularization techniques presented in [5, 16, 18, 19, 20].
@@ Acknowledgments
Harshit Bajpai would like to express gratitude to the Science and Engineering Research Board (SERB), India, for providing his PhD fellowship through grant CRG/2022/005491. Ankik Kumar Giri also extends his thanks to SERB for supporting his research under the project “Fast and efficient iterative methods for solving nonlinear inverse problems” funded by grant CRG/2022/005491.

### v3

@@ 3.2. Regularization Quality
4. Numerical Validations and Discussions
In this section, we present a series of numerical experiments to assess the practical performance of the proposed scheme, implemented in conjunction with Rule 1. The experiments are designed to illustrate the effectiveness of the method in solving ill-posed inverse problems, encompassing both linear and nonlinear cases.
@@ 4.1. Linear case
We consider the following lnear system.
((4.1)) $F_{i}(u)=\int_{a}^{b}k(t_{i},s)u(s)\,ds=y_{i},\hskip 14.22636pti=0,1,\ldots,P-1,$
which is derived from a linear integral equation on $[a,b]$, discretized by sampling at points $t_{i}\in[a,b]$ for $i=0,1,\ldots,P-1,$ where $t_{i}=a+\frac{i(b-a)}{P-1}.$ For a comprehensive discussion on such problems, see [20]. It is assumed that the kernel function $k(\cdot,\cdot)$ maintains continuity throughout the domain $a,b]\times[a,b]$. To evaluate the effectiveness of the IRSGD method (1.11), we apply it to system (4.1) with P=1000,[a,b]=[6,6],P=1000,[a,b]=[-6,6], and the kernel function given by $k(t,s)=\varphi(t-s),$ where $\varphi(t)=(1+\cos(\frac{\pi t}{3}))_{\chi_{\{|t|<3\}}}.$ The goal is to reconstruct the target solution given as
@@
In place of $y_{i}=F_{i}(u^{\dagger})$ (exact data), we consider noisy observations given by
((4.2)) $y^{\delta}_{i}=y_{i}+{\epsilon_{i}}\delta_{\text{rel}}|y_{i}|,\quad i=0,1,\ldots,P-1,$
where $\delta_{\text{rel}}$ denotes the relative noise level and $\epsilon_{i}$ represents random noise drawn from a standard Gaussian distribution $N(0,1)$. To discretize the problem, we partition the interval [6,6][-6,6] into P1P-1 equal subintervals and approximate the integrals using the trapezoidal rule. This leads to a discretized linear system of dimension $n=1000$ with the corresponding forward operator $F_{n\times n}$. We execute the method (1.11) with the numerical settings as follows:
@@
    The maximum number of iterations, denoted by $k_{\text{max}}$, is set to $1000$.
@@
    Noisy data is employed with four noise levels given by $\delta_{\text{rel}}=10^{-1},10^{-2},10^{-3}$ and $10^{-4}$.
@@ 4.1.1. Convergence behavior of $\Psi(k,y^{\delta})$
Figure 1. Convergence behavior of $\Psi(k,y^{\delta})$ for varying noise levels, illustrating the sensitivity of the proposed stopping rule to measurement perturbations.
Figure 1 clearly shows that the noise level in the input data has a significant effect on the convergence behavior of $\Psi(k,y^{\delta})$. As $\delta_{\text{rel}}$ decreases from $10^{-1}$ to $10^{-4}$, convergence becomes both faster and more stable. Additionally, it demonstrates that $\Psi(k_{*},y^{\delta})\to 0$ as $\delta\to 0$, which supports the assertion in Lemma 2.2. Furthermore, as $\delta$ decreases, the stopping index $k_{*}$ increases.
@@ 4.1.2. Reconstructed solutions
Table 1. Performance of the IRSGD method for various values of the parameter $a$ in Rule 1 and different relative noise levels $\delta_{\text{rel}}$. For each combination, the table reports the stopping index $k_{*}$ and the corresponding reconstruction error $E_{k_{*}}^{\delta}$.
@@ 4.1.3. Comparison between SGD and IRSGD
For this comparison, we set the constant $a$ in the rule (referred to as Rule 1) to $a=100$. The numerical findings are reported in Table 2.
@@
Table 2. Quantitative comparison between IRSGD and SGD methods using Rule 1 for various relative noise levels $\delta_{\text{rel}}$. The table reports the stopping index $k_{*}$, reconstruction error $E_{k_{*}}^{\delta}$, and $\Psi(k_{*},y^{\delta})$ for each case.
@@ 4.2. Nonlinear case
For each $i$ between 0 and P1P-1, let $\sigma_{i}\in S^{1}$ be a set of recording directions such that $\zeta_{i}=\zeta(\theta_{i})=(\cos\theta_{i},\sin\theta_{i})$. The objective is to recover a function $u$ supported within the bounded domain $D:=[-1,1]\times[-1,1]\subset\mathbb{R}^{2}$ from the corresponding system of equations
$F_{i}(u)=y_{i}^{\delta},\quad\text{for }i=0,1,\ldots,P-1,$
@@
$u\mapsto R_{i}(u):=\left(s\mapsto\int u(s\sigma_{i}+r\sigma_{i}^{\perp})\,dr\right).$
From [10], we know that each operator $F_{i}:H_{0}^{1}(D)\to L^{2}(I)$ with I=[1,1]I=[-1,1] is continuous and Fréchet differentiable with
$F_{i}^{\prime}(u)h=2R_{i}(u)R_{i}(h),$
for all $h\in H_{0}^{1}(D).$ Moreover, the adjoint $F_{i}^{\prime}(u)^{*}:L^{2}(I)\to H_{1}^{0}(D)$ is given by
@@
where $R_{i}^{*}:L^{2}(I)\to L^{2}(D)$ denotes the adjoint of $R_{i}$, defined by $(R_{i}^{*}g)(u)=g(u\cdot\sigma_{i}),$ II refers to the identity operator and $\Delta$ represents the Laplace operator acting on $H_{0}^{1}(D)$. Evaluating the adjoint operator entails the solution of an elliptic boundary value problem, for which a finite difference discretization is employed within a multigrid framework to obtain an efficient numerical approximation.
In this numerical experiment, the parameters were set as follows: $\delta_{\text{rel}}=10^{-2}$, $k_{\text{max}}=1000$, $\omega_{k}=\omega=5\times 10^{-3}$, and $\lambda_{k}=k^{-3}$. The true image and the corresponding initial guess are presented in Figure LABEL:initial_and_true_figures. The proposed method was applied to reconstruct the image, with the stopping criterion determined by Rule 1 for various values of the parameter $a$, as illustrated in Figure LABEL:Reconstructed.
To assess the reconstruction performance, we monitored both the relative error and the evolution of $\Psi(k,y^{\delta})$ for selected values of $a$, specifically $a=10^{2},a=10^{3}$ and $a=10^{4}$. The corresponding results are displayed in Figures LABEL:RE=1000, LABEL:RE=_10000 and LABEL:RE=_100000. Furthermore, the quality of the reconstruction was evaluated using the peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) [34], with the obtained values reported in the Table 3, alongside the stopping index and relative error for each case.
$\delta_{\text{real}}=0.01$
$a$ in Rule 1 $k_{*}$ $E_{k_{*}}^{\delta}$ $\Psi(k_{*},y^{\delta})$ PSNR SSIM
$100$ $719$ $1.62e^{-1}$ $1.29e^{+8}$ $28.11$ $0.8251$
$1000$ $693$ $1.65e^{-1}$ $2.97e^{+8}$ $27.95$ $0.8215$
$10000$ $698$ $1.70e^{-1}$ $2.04e^{+9}$ $27.70$ $0.8051$
Table 3. Performance comparison for different values of the parameter $a$ in Rule 1 for $\delta_{\text{real}}=0.01$, showing the stopping index $k_{*}$, reconstruction error $E_{k_{*}}^{\delta}$, $\Psi(k_{*},y^{\delta})$, and image quality metrics (PSNR and SSIM).
@@ 5. Conclusion
In this work, we have introduced an Iteratively Regularized Stochastic Gradient Descent (IRSGD) method for solving systems of nonlinear ill-posed inverse problems, incorporating the Hanke–Raus heuristic rule as a fully data-driven stopping criterion. A key advantage of this approach is that it operates without any prior knowledge of the noise level, thereby enhancing its applicability in practical scenarios. Our convergence analysis relies on the tangential cone condition, noise assumption and assumes Lipschitz continuity of the underlying operators. Moreover, we have established convergence results in the mean-squared norm and validated the method through comprehensive numerical experiments on both linear and nonlinear ill-posed problems. The results demonstrate that the proposed IRSGD method, augmented with a damping factor, offers improved efficiency and faster convergence compared to standard stochastic gradient approaches.
Several promising directions for future research emerge from this study. First, it would be valuable to relax Assumption 2.2 by introducing milder conditions, following the ideas in [12, 14]. Second, in line with [1, 3, 4, 26], the framework could be extended to handle non-smooth problems, thereby broadening its applicability. Finally, the development of adaptive step-size strategies could provide more practical guidelines for real-world implementation. Additionally, the heuristic stopping rule (based on a modified discrepancy principle) could be further refined to enable a deeper stochastic analysis of regularization methods such as those in [4, 22, 25, 26, 27].
@@ Acknowledgments
Harshit Bajpai acknowledges the Anusandhan National Research Foundation (ANRF), India, for supporting his Ph.D. fellowship under grant CRG/2022/005491. Ankik Kumar Giri gratefully acknowledges ANRF for funding his research through the project CRG/2022/005491.
Data Availability
Not applicable for this article.